OpenStudy (anonymous):

Im in eighth grade please help me solve this problem... Im in algebra one write an equation in slope intercept form for the line going through 2,7 and -4, 9

OpenStudy (mertsj):

Write ordered pairs like this: (2,7) (-4,9) Slope = (9-7)/(-4-2) = 2/-6 = -1/3